OpenStudy (anonymous):

i need help with a.) determining where the function is increasing or decreasing. and b.) determine where the function is concave up or concave down.

OpenStudy (anonymous):

t-ln(t)

OpenStudy (anonymous):

f(x)=t-ln(t)

OpenStudy (anonymous):

you can use the first derivative test to find part a. 2nd derivative test for part b

OpenStudy (anonymous):

well i got 1-1/t for the first derivative but im not sure if thats correct

OpenStudy (anonymous):

differentiate the function - as dockworker said

OpenStudy (anonymous):

perhaps you should post the original equation

OpenStudy (anonymous):

f(x)= t - ln(t)

OpenStudy (anonymous):

that is the original equation

OpenStudy (anonymous):

dont we need to know the range of values involved here?

OpenStudy (anonymous):

so you need to find the critical numbers (the zeros of the first derivative and also those numbers that make the first derivative undefined)

OpenStudy (anonymous):

it dosnt give me a range of values

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so t=0, t=1 are your critical numbers

OpenStudy (anonymous):

how did you get zero and one?

OpenStudy (anonymous):

test a sample from the 3 intervals these 2 critical numbers divide the domain into (-infinity,0), (0, 1), (1, infinity)

OpenStudy (anonymous):

when t=0, the derivative is undefined

OpenStudy (anonymous):

the first derivative is the slope of the tangent line. if the slope is positive in one of these intervals, the function is increasing in that interval. if negative, decreasing in that interval

OpenStudy (anonymous):

i'm sorry but i must go now. you should bump this question to the top to receive more help

OpenStudy (anonymous):

you'll find the graph of this function here http://www.wolframalpha.com/input/?i=y+%3D+t+-+ln+t

OpenStudy (anonymous):

as you can see its concave up when t > 0 second derivative = 1 / t^2 which is positive (because t^2 must be postive) and so this is a minimum whch is concave up